You choose a card at random from a standard deck of cards, replacing each card drawn, until you get a diamond. What is the probability that you will get a diamond in not more than 3 trials?


1
Expert's answer
2021-05-07T12:33:42-0400

Solution:

P(getting a diamond card in not more than 3 trials) = P(getting a diamond card in 1 trial)+P(getting a diamond card in 2 trials)+P(getting a diamond card in 3 trials)

"=\dfrac{13}{52}+\dfrac{39}{52}\times \dfrac{13}{52}+(\dfrac{39}{52})^2\times \dfrac{13}{52}
\\=\dfrac{13}{52}(1+\dfrac{39}{52}+(\dfrac{39}{52})^2)
\\=\dfrac{37}{64}"